# literature.bib

### v2

@@ 4 Stability Analysis in Discrete Time
$\mathcal{W}_{2}(\mu^{N},\tilde{\mu}^{N})\leq\mathcal{T}_{1,\lambda^{-1}}(\pi^{0},\tilde{\pi}^{0}),$
where $\mathcal{T}_{1,\lambda^{-1}}$ is the transport distance over the joint primal-dual variable defined in Section 2.3. If $\tau\sigma L^{2}<1$, then the joint variable obeys
$\mathcal{T}_{1,\lambda^{-1}}(\pi^{N},\tilde{\pi}^{N})\leq C\mathcal{T}_{1,\lambda^{-1}}(\pi^{0},\tilde{\pi}^{0}),$
@@
$\mathcal{W}_{2}(\mu^{N},\tilde{\mu}^{N})\leq\theta^{N/2}\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi^{0},\tilde{\pi}^{0}).$
@@
$\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi^{N},\tilde{\pi}^{N})\leq\tilde{C}\theta^{N/2}\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi^{0},\tilde{\pi}^{0}).$
It follows directly that in the latter case a stationary distribution $\pi^{\text{{stat}}}$ exists and the distribution $\pi^{N}$ of samples converges to $\pi^{\text{{stat}}}$ as $N\to\infty$.
In order to prove Theorems 4.1 and 4.2, we need the following auxiliary inequality, which has been used several times in similar forms for primal-dual optimization algorithms and is not specific to the sampling scheme here, see e.g. [Chambolle2011].
@@
Let $(X^{n},Y^{n})$ and $(\tilde{X}^{n},\tilde{Y}^{n})$, $n\geq 1$ be given by the coupling (4.1) with arbitrary extrapolation parameter $\theta$. Using the definition of the proximal mapping, we know that
$q^{n+1}$ $\coloneqq(Y^{n}-Y^{n+1})/\sigma+KX_{\theta}^{n}$
$p^{n+1}$ $\coloneqq(X^{n}-X^{n+1}+\sqrt{2\tau}\xi^{n+1})/\tau-K^{T}Y^{n+1},$
define analogously $\tilde{p}^{n+1},\tilde{q}^{n+1}$ for the second chain $(\tilde{X}^{n},\tilde{Y}^{n})$. Then it holds
$2\langle X^{n+1}-\tilde{X}^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle+\lVert X^{n+1}-\tilde{X}^{n+1}\rVert_{\tau^{-1}}^{2}+\lVert X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}$
$\quad+2\langle Y^{n+1}-\tilde{Y}^{n+1},q^{n+1}-\tilde{q}^{n+1}\rangle+\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle+(1-\sigma\tau\theta L^{2})\lVert Y^{n+1}-\tilde{Y}^{n+1}-Y^{n}+\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
@@
By plugging in the representation of the subgradient $p^{n+1}$ of $g$ at $X^{n+1}$ (and likewise for the second chain) given in the statement of the lemma, we can rewrite the symmetric Bregman divergence of $g$ at $(X^{n+1},\tilde{X}^{n+1})$ as
$2\langle X^{n+1}-\tilde{X}^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle$
$=2\langle X^{n+1}-\tilde{X}^{n+1},(X^{n}-X^{n+1})/\tau-K^{T}Y^{n+1}-(\tilde{X}^{n}-\tilde{X}^{n+1})/\tau+K^{T}\tilde{Y}^{n+1}\rangle$
$=-\frac{2}{\tau}\langle X^{n+1}-\tilde{X}^{n+1},X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}\rangle-2\langle X^{n+1}-\tilde{X}^{n+1},K^{T}(Y^{n+1}-\tilde{Y}^{n+1})\rangle$
$=\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}-\lVert X^{n+1}-\tilde{X}^{n+1}\rVert_{\tau^{-1}}^{2}-\lVert X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}$
$\quad-2\langle K(X^{n+1}-\tilde{X}^{n+1}),Y^{n+1}-\tilde{Y}^{n+1}\rangle,$
@@
$2\langle Y^{n+1}-\tilde{Y}^{n+1},q^{n+1}-\tilde{q}^{n+1}\rangle$
$=2\langle Y^{n+1}-\tilde{Y}^{n+1},(Y^{n}-Y^{n+1})/\sigma+KX_{\theta}^{n}-(\tilde{Y}^{n}-\tilde{Y}^{n+1})/\sigma-K\tilde{X}_{\theta}^{n}\rangle$
$=-\frac{2}{\sigma}\langle Y^{n+1}-\tilde{Y}^{n+1},Y^{n+1}-\tilde{Y}^{n+1}-Y^{n}+\tilde{Y}^{n}\rangle+2\langle Y^{n+1}-\tilde{Y}^{n+1},K(X_{\theta}^{n}-\tilde{X}_{\theta}^{n})\rangle$
$=\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}-\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert_{\sigma^{-1}}^{2}-\lVert Y^{n+1}-\tilde{Y}^{n+1}-Y^{n}+\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\langle K(X_{\theta}^{n}-\tilde{X}_{\theta}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle.$
In order to shorten the expressions, we introduce the shorthand notation $U^{n}:=X^{n}-\tilde{X}^{n}$, $U_{\theta}^{n}=X_{\theta}^{n}-\tilde{X}_{\theta}^{n}$ and $V^{n}:=Y^{n}-\tilde{Y}^{n}$. Adding up the two above equalities gives
@@
$\quad+\theta\lVert U^{n}-U^{n-1}\rVert_{\tau^{-1}}^{2}+\tau\sigma\theta L^{2}\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}.$
The inequality follows from applying Cauchy-Schwarz and the inequality $ab\leq a^{2}/2c+cb^{2}/2$ for any $c>0$ (Young’s inequality, with $a=\lVert K(U^{n}-U^{n-1})\rVert$ and $b=\lVert V^{n+1}-V^{n}\rVert$ and $c=\tau L^{2}$ here) to the last inner product in the expression before. Inserting this bound into (4) gives precisely the statement of the Lemma. ∎
@@
$\langle X^{n+1}-\tilde{X}^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle$ $\geq 0,$
$\langle Y^{n+1}-\tilde{Y}^{n+1},q^{n+1}-\tilde{q}^{n+1}\rangle$ $\geq 0.$
@@
Iterating this inequality over $n=0,\dots,N-1$ steps and using $X^{-1}=X^{0}$, $\tilde{X}^{-1}=X^{0}$ gives
@@
$\quad+2\langle K(X^{N}-\tilde{X}^{N}-X^{N-1}+\tilde{X}^{N-1}),Y^{N}-\tilde{Y}^{N}\rangle+\left(1-\tau\sigma L^{2}\right)\sum_{n=0}^{N-1}\lVert Y^{n+1}-\tilde{Y}^{n+1}-Y^{n}+\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
@@
Using Cauchy-Schwarz and Young’s inequality on the remaining inner product and non-negativity of the sum of norms implies
@@
Taking expectations on both sides, by definition of the Wasserstein distance, the left side is lower bounded by $\tau^{-1}\mathcal{W}_{2}^{2}(\mu^{N},\tilde{\mu}^{N})$. Since the resulting inequality holds for all couplings of initializations $(X^{0},Y^{0})$ and $(\tilde{X}^{0},\tilde{Y}^{0})$, taking the infimum over all couplings on the right side gives $\tau^{-1}\mathcal{T}^{2}_{1,\lambda^{-1}}(\pi^{0},\tilde{\pi}^{0})$, proving the first inequality. The second inequality in the statement of Theorem 4.1 results from lower bounding the left side not by the Wasserstein distance in the primal variable, but instead by $C^{-2}\tau^{-1}\mathcal{T}^{2}_{1,\lambda^{-1}}(\pi^{N},\tilde{\pi}^{N})$. ∎
@@
$\langle X^{n+1}-\tilde{X}^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle$ $\geq\omega_{g}\lVert X^{n+1}-\tilde{X}^{n+1}\rVert^{2},$
$\langle Y^{n+1}-\tilde{Y}^{n+1},q^{n+1}-\tilde{q}^{n+1}\rangle$ $\geq\omega_{f^{\ast}}\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert^{2}.$
@@
$\quad+2\langle K(X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle+(1-\sigma\tau\theta L^{2})\lVert Y^{n+1}-\tilde{Y}^{n+1}-Y^{n}+\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
@@
$\quad 2\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle.$
@@
$\theta\Delta^{n}$ $\geq\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad 2\theta\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle$
$\geq\Delta^{n+1}+(1-\sigma\tau\theta L^{2})\lVert Y^{n+1}-\tilde{Y}^{n+1}-Y^{n}+\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\geq\Delta^{n+1}$
@@
$\Delta^{0}$ $\geq\theta^{-N}\Delta^{N}.$

### v3

@@ 4.1 Stability of Primal-Dual Langevin Sampling
$\mathcal{W}_{2}(\mu_{N},\tilde{\mu}_{N})\leq\mathcal{T}_{1,\lambda^{-1}}(\pi_{0},\tilde{\pi}_{0}),$
where $\mathcal{T}_{1,\lambda^{-1}}$ is the transport distance over the joint primal-dual variable defined in (3.1). If $\tau\sigma L^{2}<1$, then the joint variable obeys
$\mathcal{T}_{1,\lambda^{-1}}(\pi_{N},\tilde{\pi}_{N})\leq C\mathcal{T}_{1,\lambda^{-1}}(\pi_{0},\tilde{\pi}_{0}),$
@@
$\mathcal{W}_{2}(\mu_{N},\tilde{\mu}_{N})\leq\theta^{N/2}\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi_{0},\tilde{\pi}_{0}).$
@@
$\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi_{N},\tilde{\pi}_{N})\leq\tilde{C}\theta^{N/2}\mathcal{T}_{1,\tilde{\lambda}^{-1}}(\pi_{0},\tilde{\pi}_{0}).$
It follows directly that in the latter case a stationary distribution $\pi_{\lambda,\tau}$ exists and the distribution $\pi_{N}$ of samples converges to $\pi_{\lambda,\tau}$ as $N\to\infty$.
In order to prove Theorems 4.1 and 4.2, we need the following auxiliary inequality, which has been used several times in similar forms for primal-dual optimization algorithms, see e.g. [Chambolle2011]. It still holds in the present sampling setting, since considering the difference of two coupled processes makes the stochastic terms $\xi^{n}$ cancel out:
@@
Let $(X^{n},Y^{n})$ and $(\tilde{X}^{n},\tilde{Y}^{n})$, $n\geq 0$ be given by the coupling (4.1) with extrapolation parameter $\theta$ chosen such that $\theta\sigma\tau L^{2}\leq 1$. By definition of the proximal mapping, we have for $n\geq 0$:
$\begin{array}[]{l}p^{n+1}\coloneqq(X^{n}-X^{n+1}+\sqrt{2\tau}\xi^{n+1})/\tau-K^{T}Y^{n+1},\in\partial g(X^{n+1}-\sqrt{2\tau}\xi^{n+1})\vspace{0.5em}\\ q^{n+1}\coloneqq(Y^{n}-Y^{n+1})/\sigma+KX_{\theta}^{n}\in\partial f^{\ast}(Y^{n+1}).\end{array}$ (4.2)
Analogously, we define $\tilde{p}^{n+1},\tilde{q}^{n+1}$ for the second chain $(\tilde{X}^{n},\tilde{Y}^{n})$. Then it holds
$2\left(D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})+D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})\right)$
$\quad+\lVert X^{n+1}-\tilde{X}^{n+1}\rVert_{\tau^{-1}}^{2}+\lVert X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\langle K(X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle$
@@
In order to shorten the expressions, throughout the proof we abbreviate $U^{n}:=X^{n}-\tilde{X}^{n}$, $U_{\theta}^{n}=X_{\theta}^{n}-\tilde{X}_{\theta}^{n}$ and $V^{n}:=Y^{n}-\tilde{Y}^{n}$.
By plugging in the representation of the subgradient $p^{n+1}$ of $g$ at $X^{n+1}-\sqrt{2\tau}\xi^{n+1}$ (and likewise $\tilde{p}^{n+1}$) given in the statement of the lemma, we can rewrite a symmetric Bregman divergence as
$2D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})$
$\quad=2\langle U^{n+1},p^{n+1}-\tilde{p}^{n+1}\rangle$
$\quad=2\langle U^{n+1},(X^{n}-X^{n+1})/\tau-K^{T}Y^{n+1}-(\tilde{X}^{n}-\tilde{X}^{n+1})/\tau+K^{T}\tilde{Y}^{n+1}\rangle$
$\quad=\frac{2}{\tau}\langle U^{n+1},U^{n}-U^{n+1}\rangle-2\langle U^{n+1},K^{T}V^{n+1}\rangle$
$\quad=\lVert U^{n}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}\rVert_{\tau^{-1}}^{2}-\lVert U^{n+1}-U^{n}\rVert_{\tau^{-1}}^{2}-2\langle KU^{n+1},V^{n+1}\rangle,$
@@
$2D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})$ $=2\langle V^{n+1},(Y^{n}-Y^{n+1})/\sigma+KX_{\theta}^{n}-(\tilde{Y}^{n}-\tilde{Y}^{n+1})/\sigma-K\tilde{X}_{\theta}^{n}\rangle$
$=\frac{2}{\sigma}\langle V^{n+1},V^{n}-V^{n+1}\rangle+2\langle V^{n+1},KU_{\theta}^{n}\rangle$
$=\lVert V^{n}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}\rVert_{\sigma^{-1}}^{2}-\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}+2\langle KU_{\theta}^{n},V^{n+1}\rangle.$
Adding up the two equalities above gives
@@
$\quad+\theta\lVert U^{n}-U^{n-1}\rVert_{\tau^{-1}}^{2}+\tau\sigma\theta L^{2}\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}.$ (4.4)
Inserting this into (4.1) and using $(1-\sigma\tau\theta L^{2})\lVert V^{n+1}-V^{n}\rVert_{\sigma^{-1}}^{2}\geq 0$ gives the statement of the Lemma. ∎
@@
$D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})\geq 0,$
$D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})\geq 0.$
@@
Iterating this inequality over $n=0,\dots,N-1$ steps and using $X^{-1}=X^{0}$, $\tilde{X}^{-1}=\tilde{X}^{0}$ gives
@@
$\quad+2\langle K(X^{N}-\tilde{X}^{N}-X^{N-1}+\tilde{X}^{N-1}),Y^{N}-\tilde{Y}^{N}\rangle$
@@
Using the bound $2\langle a,b\rangle\geq-c^{-1}\lVert a\rVert^{2}-c\lVert b\rVert^{2}$ with $c=\tau L^{2}$ on the remaining inner product implies
@@
Taking expectations on both sides, by definition of the Wasserstein distance, the left side is lower bounded by $\tau^{-1}\mathcal{W}_{2}^{2}(\mu_{N},\tilde{\mu}_{N})$. Since the resulting inequality holds for all couplings of initializations $(X^{0},Y^{0})$ and $(\tilde{X}^{0},\tilde{Y}^{0})$, taking the infimum over all couplings on the right side gives $\tau^{-1}\mathcal{T}^{2}_{1,\lambda^{-1}}(\pi_{0},\tilde{\pi}_{0})$, proving the first inequality. The second inequality in the statement of Theorem 4.1 results from lower bounding the left side not by the Wasserstein distance in the primal variable, but instead by $C^{-2}\tau^{-1}\mathcal{T}^{2}_{1,\lambda^{-1}}(\pi_{N},\tilde{\pi}_{N})$. ∎
@@
$D_{g,\text{sym}}^{p^{n+1},\tilde{p}^{n+1}}(X^{n+1}-\sqrt{2\tau}\xi^{n+1},\tilde{X}^{n+1}-\sqrt{2\tau}\xi^{n+1})\geq\omega_{g}\lVert X^{n+1}-\tilde{X}^{n+1}\rVert^{2},$
$D_{f^{\ast},\text{sym}}^{q^{n+1},\tilde{q}^{n+1}}(Y^{n+1},\tilde{Y}^{n+1})\geq\omega_{f^{\ast}}\lVert Y^{n+1}-\tilde{Y}^{n+1}\rVert^{2}.$
@@
$\quad+2\langle K(X^{n+1}-\tilde{X}^{n+1}-X^{n}+\tilde{X}^{n}),Y^{n+1}-\tilde{Y}^{n+1}\rangle$
@@
$\Delta^{n+1}$ $\leq\lVert X^{n}-\tilde{X}^{n}\rVert_{\tau^{-1}}^{2}+\theta\lVert X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}\rVert_{\tau^{-1}}^{2}+\lVert Y^{n}-\tilde{Y}^{n}\rVert_{\sigma^{-1}}^{2}$
$\quad+2\theta\langle K(X^{n}-\tilde{X}^{n}-X^{n-1}+\tilde{X}^{n-1}),Y^{n}-\tilde{Y}^{n}\rangle$
$\leq\theta\Delta^{n}$
@@
$\Delta^{N}$ $\leq\theta^{N}\Delta^{0}.$